solve the equation
3(z-9) -z(z-9)=0
Remove the brackets :
3z-27-z2+9z=0
add so:
z2-9z+27=0
factorizing
(z-9)(z-3)=0
z-9=0 and z-3=0
ie:
z= 9 and z=3

Answer by IBRAHIM92(5)

solve the equation
3(z-9) -z(z-9)=0
Remove the brackets :
3z-27-z2+9z=0
add so:
z2-9z+27=0
factorizing
(z-9)(z-3)=0
z-9=0 and z-3=0
ie:
z= 9 and z=3